# MetaSlot: Break Through the Fixed Number of Slots in Object-Centric Learning

### v1

@@ Appendix D Experimental Details
To ensure a fair comparison, we re-implemented all baseline models from scratch rather than relying on publicly reported results. Throughout all experiments, we kept data augmentation strategies, the visual feed-forward module (VFM) in the OCL encoder—based on DINOv2 ViT-s/14 [50]—and all training hyperparameters identical to those reported in the original papers. Furthermore, we replaced each model’s original variational autoencoder (VAE) component with a large-scale pre-trained TAESD module [75], which is based on Stable Diffusion.
@@
All models—including both the MetaSlot-augmented variants and their respective baselines—were trained using the Adam optimizer [69] on a single NVIDIA V100 GPU with 16-bit mixed precision. Each run consisted of 50000 training steps, with a batch size of 32 and four data-loading workers. We set the initial learning rate to $2\times 10^{-4}$ and maintained it throughout training. For the MetaSlot module, the codebook size was fixed at 512, the feature map resolution at $256\times 256$, and the embedding dimension at 256. The number of slots for each model remained consistent with its original baseline configuration. To reduce the impact of randomness, all experiments were repeated with three different random seeds, and we report the mean results.
Pre-trained VQ-VAE Configuration for SLATE and Slot Diffusion.
@@
For the SLATE and Slot Diffusion baselines, we adhered to the standard VQ-VAE implementation based on ResNet-18. Specifically, we used a codebook size of 4096 and an embedding dimension of 256. Pre-training was conducted for 30000 steps with a batch size of 64 shared between the text and vision branches, four data-loading workers, and an initial learning rate of $2\times 10^{-3}$. The feature map resolution remained $256\times 256$. As before, each configuration was evaluated over three independent runs, and the reported metrics represent the averaged performance across these trials.

### v2

@@ Appendix C Experimental Details
To ensure a fair comparison, we re‐implemented all baseline models from scratch rather than relying on publicly reported results. Throughout all experiments, we kept data augmentation strategies, the visual feed‐forward module (VFM) in the OCL encoder—based on DINOv2 ViT-s/14 [50]—and all training hyperparameters identical to those reported in the original papers. Furthermore, we replaced each model’s original variational autoencoder (VAE) component with a large‐scale pre‐trained TAESD module [78], which is based on Stable Diffusion.
@@
All models—including both the MetaSlot‐augmented variants and their respective baselines—were trained using the Adam optimizer [71] on a single NVIDIA V100 GPU with 16‑bit mixed precision. Each run consisted of 50000 training steps, with a batch size of 32 and four data‑loading workers. We set the initial learning rate to $2\times 10^{-4}$ and maintained it throughout training. For the MetaSlot module, the codebook size was fixed at 512, the feature map resolution at $256\times 256$, and the embedding dimension at 256. The number of slots for each model remained consistent with its original baseline configuration. To reduce the impact of randomness, all experiments were repeated with three different random seeds, and we report the mean results.
Pre‐trained VQ‑VAE Configuration for SLATE and Slot Diffusion.
@@
For the SLATE and Slot Diffusion baselines, we adhered to the standard VQ‑VAE implementation based on ResNet‑18. Specifically, we used a codebook size of 4096 and an embedding dimension of 256. Pre‑training was conducted for 30000 steps with a batch size of 64 shared between the text and vision branches, four data‑loading workers, and an initial learning rate of $2\times 10^{-3}$. The feature map resolution remained $256\times 256$. As before, each configuration was evaluated over three independent runs, and the reported metrics represent the averaged performance across these trials.
Appendix D Supplementary Ablation Experiments
Impact of Codebook Size
As shown in table 5, we investigated the impact of different codebook sizes (256, 512, and 1024) on model performance. We observed that codebook size has a limited effect on performance; however, a larger prototype set allows for finer distinctions between slots that share semantic concepts but differ in spatial location. For example, as shown in Fig. 3, slot #445 corresponds to the concept of “person” located on the right side of the image, while slot #337 represents the same concept but appears at the center of the scene.
Impact of Slot Number
As shown in table 6, we examined how the number of slots affects final performance. Since the codebook and the aggregator module are jointly optimized, the quality of the aggregator—particularly in early training stages—can significantly influence codebook convergence. We found that the empirically optimal slot count aligns well with long-standing choices commonly adopted in the community.
Table 5: Results of MetaSlot with varying codebook sizes on MS COCO.
DINOSAUR on MS COCO
ARI FG-ARI MBO mIoU
MetaSlot, codebook_size=256 26.7 38.1 29.4 27.5
MetaSlot, codebook_size=512 22.4 40.3 29.5 27.9
MetaSlot, codebook_size=1024 20.0 40.4 29.2 27.3
Table 6: Results of MetaSlot with varying slot numbers on MS COCO.
DINOSAUR on MS COCO
ARI FG-ARI MBO mIoU
MetaSlot, slot_num=5 25.5 37.8 29.3 27.4
MetaSlot, slot_num=7 22.4 40.3 29.5 27.9
MetaSlot, slot_num=11 19.8 38.7 28.4 26.9
MetaSlot, slot_num=15 15.9 36.0 27.2 26.0
Impact of Architectural Components
As shown in table 8, we conducted additional experiments to gain deeper insight into the working mechanism of MetaSlot. Specifically, we tested two variants: (1) removing the prototype guidance and instead initializing slots in the second stage by sampling from separate Gaussian distributions, as done in the first stage; and (2) disabling the reactivation mechanism for stale (dead) prototypes during the codebook update process. The results confirm the effectiveness of our module design and are consistent with the theoretical assumptions proposed in the paper.
Furthermore, prior studies in object-centric learning have consistently shown that simply increasing the number of Slot Attention iterations does not lead to meaningful performance gains. To validate this observation, we performed an additional ablation study on the BO-QSA model by increasing its iteration count to 6, matching that of MetaSlot. As shown in the results, increasing the iteration count alone does not improve BO-QSA’s performance, further highlighting the importance of our design choices beyond iteration depth.
Table 7: Supplementary ablation on architectural components. COCO #slot=7 FG-ARI mBO mIoU MetaSlot w/o proto 40.2 29.1 27.6 MetaSlot w/o prune 40.0 29.1 27.5 MetaSlot 40.3 29.5 27.9 BO-QSA,iter_num=6 37.9 27.7 26.3    Table 8: Comparison with SlotContrast on the MOVi-C dataset. Method FG-ARI mBO SlotContrast 62.4 30.6 MetaSlot 63.9 35.0
Appendix E Supplementary Comparative Experiments
Comparative Evaluation against SlotContrast
As shown in table 10, we further include a comparison with the SlotContrast[76] model on the MOVi-C dataset. In these experiments, MetaSlot does not employ SlotContrast losses or any temporal-specific enhancements. Due to constraints in time and computational resources, we used a batch size of 8 (vs. 64 in SlotContrast), 24 frames per sample, and trained for up to 20,000 steps (vs. 100,000 in SlotContrast). Despite these limitations, MetaSlot still demonstrates notable performance advantages.
It is worth noting that SlotContrast’s contrastive loss was originally designed for a fixed number of slots, and extending it to handle variable slot counts would require additional effort. Nevertheless, as discussed earlier, even without incorporating SlotContrast’s core contribution—the contrastive loss—MetaSlot achieves comparable or even superior performance in object discovery tasks. We believe that integrating SlotContrast’s contrastive learning objective in the first stage with MetaSlot’s dynamic aggregation in the second stage could further enhance performance, particularly for temporal object discovery. However, this would necessitate architectural modifications that are beyond the scope of the current work and are left for future research.
Table 9: Comparison with SlotContrast on the MOVi-C dataset. Method FG-ARI mBO SlotContrast 62.4 30.6 MetaSlot 63.9 35.0    Table 10: Comparison with AdaSlot on the MS COCO dataset. Method FG-ARI mBO AdaSlot 39.0 27.36 BO-QSA 37.64 27.07 MetaSlot 40.13 27.45
Comparative Evaluation against AdaSlot
As shown in table 10, we further report the results of the MetaSlot model using ViT-B/16 under the same training settings as in the AdaSlot[53] paper on MS COCO dataset. Due to limitations in time and computational resources, MetaSlot was trained for only 40,000 steps (compared to 500,000 steps in AdaSlot), yet it already surpasses the best performance reported in the AdaSlot paper.
We speculate that this is partly because, when the number of slots varies, the model can no longer apply separate Gaussian initialization to each slot individually.
Additional Evaluation of VQ-based Object-Centric Methods
We further evaluated several object-centric learning methods that employ vector quantization. In SysBinder[65] and NLoTM[66], the quantization mechanisms operate at the attribute level within each slot. However, this design does not necessarily lead to improved representation quality in unsupervised settings. In synthetic datasets, object attributes such as color, size, shape, and material are relatively fixed and can often be described using fewer than ten shared labels. In contrast, real-world datasets such as COCO and VOC contain objects with far more diverse and non-shared attributes. For instance, the object person cannot be meaningfully described with the same attribute set as the object mouse. This discrepancy likely explains why SysBinder and NLoTM were not originally evaluated on real-world benchmarks.
Furthermore, the initial intuition behind MetaSlot is inspired by Platonic philosophy. In Plato’s theory of Forms, sensible particulars are intelligible only insofar as they “participate” (metechē) in a transcendent Form (Phaedo 100c–d). Similarly, Aristotle maintains in the Categories that accidents belong to substances only as predicates of discourse, not as essential ingredients shared by all beings (Categories 2a11–19).
Our empirical evaluation on COCO and ClevrTex shows that MetaSlot substantially outperforms both SysBinder and NLoTM. Although SysBinder achieves a slight advantage in FG-ARI on ClevrTex, it underperforms the DINOSAUR baseline on other key metrics such as ARI, MBO, and mIoU. Moreover, the SysBinder paper only reports FG-ARI and does not include ARI, MBO, or mIoU—metrics widely regarded as standard for evaluating object discovery.
Table 11: Comparative results of SysBinder, NLoTM, DINOSAUR, and MetaSlot on MS COCO (unsupervised object discovery). MS COCO ARI FG-ARI MBO mIoU SysBinder 16.8 37.7 27.2 26.0 NLoTM 57.8 14.7 23.6 17.8 DINOSAUR 18.2 35.0 28.3 26.9 MetaSlot 22.4 40.3 29.5 27.9    Table 12: Comparative results of SysBinder, NLoTM, DINOSAUR, and MetaSlot on ClevrTex. ClevrTex ARI FG-ARI MBO mIoU SysBinder 21.7 91.4 46.8 46.5 NLoTM 21.5 88.7 44.1 43.1 DINOSAUR 50.7 89.4 53.3 52.8 MetaSlot 64.6 89.6 55.2 54.5